The specific conductance of a saturated solution of AgCl is 1.55*10-6

ohm-1 cm-1

and 

that of water is 4.3*10-7

ohm-1 cm-1

. In a field of 1 volt per cm. mobility of chloride and 

silver ions are 5.6*10-4 and 6.8*10-4 cm/sec. calculate solubility and solubility product 

of AgCl.


1
Expert's answer
2021-05-17T04:08:48-0400

"K_{AgCl} = K_{AgCl}(solution)-K_{H_2O}"


"= 1.55 \times 10^{-6} - 4.3 \times 10^{-7}"

"= 1.12 \times 10^{-6}"

Therefore,


"S = \dfrac{K \times 1000 }{137.2}"

"S = \dfrac{1.12 \times 10^{-6} \times 1000}{137.2}"


"S = 0.00816 \times 10^{-3}"